# NON-FACTORIZABLE RIBBON HOPF ALGEBRAS

### v1

@@ 1 Introduction
Unimodular, ribbon categories introduced by Turaev in [Tur94] constitute an important topic in quantum algebra, chiefly for their many applications in quantum topology. Since the 1990s, multiple constructions of 3-manifold and 4-manifold invariants have been constructed using these essential ingredients [Tur94], [KL01], [CY93]. Ribbon Hopf algebras are a natural source of such categories through their representation theory. A ribbon Hopf algebra $H$ is a quasitriangular Hopf algebra (i.e. it comes with the extra data of an invertible element $R$ of $H\otimes H$ called the R-matrix) together with a central element of $H$ called the ribbon element. Both $R$ and $v$ are required to satisfy some equations, see Section 2.
In particular, one may define a $3$-manifold invariant from any so-called factorizable unimodular ribbon Hopf algebra. Factorizability is a non-degeneracy condition related to the monodromy matrix $M:=R_{21}R$ of the algebra, where $R_{21}$ is the R-matrix with permuted entries (see Prop. 2.10). Recall also that a Hopf algebra $H$ is unimodular if it admits an integral $\lambda\in H^{*}$ and a cointegral $\Lambda\in H$ satisfying some properties, as detailed in Def. 2.5.
@@
At the topological level, non-factorizability is directly related to the ability of the invariants from [BdR23] to differentiate between $1$-handles and $2$-handles, which in turn is related to the ability of the invariant to see beyond the boundary of the $4$-dimensional $2$-handlebody. Non (co)semi-simplicity is related to the sensitivity of the invariant to a stabilization operation on handlebodies. This is detailed in Section 5.
@@ 1.1 Main results
Hence, in an attempt to avoid such “degeneration to the boundary”, in this paper, we introduce the stricter notion of strong non-factorizability. We demand that the result of applying the integral to either side of the monodromy is zero, and hence cannot be related to the cointegral by any invertible element of the algebra:
$(\lambda\otimes\mathrm{Id})M=0.$
@@
First we discuss a family of Hopf algebras that was introduced in [BDG00] and analysed from the angle of quasitriangularity and ribbonness in [Nen04]. We will call these Hopf algebras Nenciu algebras. Let $\mathbf{m}\in\mathbb{Z}^{s}$ be a row tuple of integers of length $s\in\mathbb{Z}$, $t\in\mathbb{Z}$ and $\boldsymbol{\mathsf{d}},\boldsymbol{\mathsf{u}}$ be $t\times s$ matrices with integer entries. A Nenciu algebra is generated by grouplike generators $K_{a}$, $a=1,\dots,s$ such that $K^{m_{a}}_{a}=\mathbf{1}$, commuting with each other, as well as skew-primitive nilpotent generators $X_{k}$, $k=1,\dots,t$ such that $X^{2}_{k}=0$. The relations between the generators are prescribed by the matrices $\boldsymbol{\mathsf{d}},\boldsymbol{\mathsf{u}}$ and are diagonal, that is commuting any two elements results only in a constant in front of the swapped product. We denote the resulting algebra by $H(\mathbf{m},t,\boldsymbol{\mathsf{d}},\boldsymbol{\mathsf{u}})$, and a full definition is provided in Definition 3.1. In [Nen04], an array of previously existing quasitriangular Hopf algebra constructions, including the family $\operatorname{SF}_{2n}$, were unified, and sufficient conditions for such an algebra to be ribbon where determined, but no new ones were given. In particular, the questions of unimodularity and factorizability were not adressed.
In this paper, we exhibit new examples. Moreover we determine sufficient conditions for unimodularity, and for a number of instances we establish when non-factorizable and strongly non-factorizable ribbon structures are admitted, relying in particular on a theorem of Radford [Rad94], reproduced below as Theorem 2.9. We also study another property: the algebra is called anomalous if $\lambda(v)\neq 1$ and anomaly-free otherwise (see Definition 2.20). Topologically, as explained in [BdR23], when the algebra is both factorizable and anomaly-free, the $4$-dimensional $2$-handlebody invariant associated to the algebra is actually an invariant of the boundary of the handlebody. For completeness, we check that anomaly-freeness is not satisifed in the examples where the corresponding quasitriangular structure is strongly non-factorizable. Combining the results of [Nen04] with our owns, we have the following.
@@
However, we again observed that no true 4-dimensional information seems to be carried by the $4$-dimensional $2$-handlebody invariants given by this family of Hopf algebras. More precisely, computations for multiple examples suggest that the invariant might be given by the product of the 2-dimensional invariant of the spine of the handlebody obtained from $H(\mathbf{m},t,\boldsymbol{\mathsf{d}},\mathbf{u})$ and 3-dimensional invariant of its boundary associated to $u_{q}\mathfrak{sl}_{2}$. Thus, we leave it as Conjecture 5.11.
@@ 1.2 Structure of the paper
The paper is organised as follows: in Section 2 we introduce all the relevant definitions pertaining to unimodular ribbon Hopf algebras and the notation used to describe the Nenciu algebras. In Section 3 the Nenciu algebras are introduced together with the conditions for a given one to admit unimodular, ribbon, and strongly non-factorizable structures. In Section 4, we introduce the second family of examples and once again give conditions for strong non-factorizability. In Section 5 we recall some topological background, including the Kerler-Lyubashenkno invariants of 4-dimensional 2-handlebodies, and discuss the efficacy of the invariants provided by our examples. Finally in Appendix A we list all the examples of Hopf algebras found throughout the paper, and in Appendix B we collect some proofs that are too long to be presented in the main body.
Acknowledgements: The authors would like to thank Ivelina Bobtcheva, Marco de Renzi, Azat Gainutdinov and Cris Negron for early comments on the paper and fruitful discussions, as well as Riccardo Piergallini for making his Mathematica computer program available, and Anna Beliakova for her help and introducing them to the topic. Both authors were supported by Simons Collaboration on New Structures in Low-Dimensional Topology and Grant 200020_207374 of the Swiss National Science Foundation.
@@ 2.3 Quasitriangular and ribbon structures
A unimodular, quasitriangular Hopf algebra $(H,R)$ with a two-sided integral $\lambda\in H^{*}$ is called strongly non-factorizable if any of the two equivalent conditions is fulfilled
$\lambda(M^{\prime})M^{\prime\prime}=0,$
@@
$M^{\prime}\lambda(M^{\prime\prime})=0.$
@@ 3 The Nenciu construction
The Nenciu construction of [Nen04] produces finite-dimensional ribbon Hopf algebras generated by grouplike generators collected in a tuple $\mathbf{K}=(K_{1},\dots,K_{s})$ and nilpotent generators collected in a tuple $\mathbf{X}=(X_{1},\dots,X_{t})$, such that the group subalgebra $\mathbb{C}[\mathbf{K}]$ is abelian, the nilpotent generators are skew-primitive with respect to $\mathbb{C}[\mathbf{K}]$ and all commutation relations between generators are diagonal. The choice of the generators and relations is strongly constrained, and sufficient conditions for the result to be a ribbon Hopf algebra are determined in [Nen04]. We will see that strongly non-factorizable examples of this construction exist.
@@ 3.1 Hopf algebra structure
$\boldsymbol{\xi}^{\boldsymbol{\mathsf{d}}_{k}\cdot\boldsymbol{\mathsf{u}}_{k}}=-1.$
@@ 3.4 Quasitriangular structure
The following is a Corollary of Prop. 3.16.
@@
    since $\mathbf{d_{k}}\boldsymbol{\mathsf{z}}=-\boldsymbol{\mathsf{u}}_{k}$. In the latter expression combining the two sums yields
@@
    Each piece of the rightmost sum corresponds in the former tensor factor to a monomial in $X_{k}$ generators, $\mathbf{X}^{\mathbf{r}}$, and in the latter tensor factor its image under the composition of the duality (which, recall, is only a linear map), and the bialgebra map so the element $f((\mathbf{X}^{\mathbf{r}})^{*})\in H^{cop}$. The parameterization coefficients $\alpha_{kl}$ control whether a term appears in the sum, or equivalently in the image of $f$, or not, and with which coefficient.
@@ 3.6 Examples revisited
Here we stick to the notation for the nilpotent generators from the previous examples emphasizing the types. In particular, let $Z^{+}_{l}$ and $Z^{-}_{l}$ be nilpotent generators in $H(\mathbf{m},t,\boldsymbol{\mathsf{d}},\boldsymbol{\mathsf{u}})$ of opposite type, with neighbouring indices, that is $Z^{+}_{l}=X_{k}$ and $Z^{-}_{l}=X_{k+1}$ for some $k$. Let moreover $\{Z_{l}^{\pm},X_{k}\}=0$ for all $X_{k}$, including themselves. We changed the notation in the expression above and we denoted $\alpha_{k,k+1}=-\alpha_{k+1,k}:=\alpha_{l}$. We collect $\alpha_{l}$ in the tuple $\boldsymbol{\alpha}$.
@@
Finally, from the definitions of the integral and the R-matrix, we observe that all nilpotent generators except $Z^{\pm}_{l}$ (or all in Example 3.10) are absent from the respective R-matrices. There is no relation available that introduces them, as the commutations are necessarily diagonal, thus they are absent from the monodromy matrix as well. Since the integral is non-zero only on the linear combinations of monomials containing the top element $T=\mathbf{X}^{(1,\dots,1)}$, and since $T$ contains all the nilpotent generators, we retrieve the result. ∎
@@ 4.1 The small quantum group $u_{q}\mathfrak{sl}_{2}$
We use this section to set the conventions for the small quantum group. We follow mostly the conventions of [CP95][Section 9.1] and [EGNO16][Section 5.7], used also in [BdR23]. Let $r^{\prime}=r/2$ and $r^{\prime\prime}=r/4$ throughout.
@@ 4.2 The semi-direct biproduct structure
We emphasize that all tensor products between $U$ and $H$ in Definition 4.4 are taken in the category $\text{{Vect}}_{\mathbb{C}}$, which carries the trivial braiding. In particular, this is a different setting than the Radford biproduct which is taken in a braided category of Yetter-Drinfeld modules over the acting algebra, see [AS06].
@@ 4.3 Non-factorizable biproducts with $u_{q}\mathfrak{sl}_{2}$
This follows directly from the fact that both $U$ and $H$ have monomial bases, by the Poincaré-Birkhoff-Witt theorem (see for instance [Kas12]) and Proposition 3.6, respectively. Indeed, all newly introduced relations between the two bases are diagonal by Theorem 4.7. Hence by the same argument as in the proof of Proposition 3.6, we can establish the existence of the monomial basis. ∎
@@
    and the corresponding pivotal element $g=L$ and the Drinfeld element
    $u_{H}:=L\exp\left(-2\sum^{t_{2}}_{k=1}\alpha_{l}Z^{+}_{l}Z^{-}_{l}\right).$
@@
By the reasoning of Proposition 3.28, if $H=H(\mathbf{m},t,\boldsymbol{\mathsf{d}},\boldsymbol{\mathsf{u}})$ is strongly non-factorizable, the top element $\mathbf{X}^{(1,\dots,1)}$ is missing from the monodromy matrix $M_{H}\in H\otimes H$. Since all new relations are also diagonal, it will be missing from the new monodromy matrix $M\in(U\ltimes H)\otimes(U\ltimes H)$, so by the same argument the result follows. ∎
@@ 4.4 Examples
$\Delta(K_{1})=K_{1}\otimes K_{1},\;\;\Delta(K_{2})=K_{2}\otimes K_{2},\;\;\Delta(X^{\pm})=\mathbf{1}\otimes X^{\pm}+X^{\pm}\otimes\bar{L}^{\pm 1}.$
@@
    $\Delta(E)=\mathbf{1}\otimes E+E\otimes K.$
@@
Thus we see the top monomial picked up by the integral of $U\ltimes H$ is absent from the monodromy matrix, as in Proposition 3.28. We conclude the proof with the remark that we need not ask about the fate of generators $K_{a}$ under the map $f$, as this does not affect the strong non-factorizability condition. ∎
@@ 5.2 Invariants of handlebodies up to $2$-equivalence
A very important feature of a unimodular ribbon Hopf algebra $H$ is that it can be ”transmuted” into a $\operatorname{BP}$-Hopf algebra $\underline{H}$ in the category $H$-$\operatorname{mod}$ (as a special case of [BdR23, Prop. 7.3]), where $\underline{H}$ is $H$ as an $H$-module for the left adjoint action from Proposition 3.8. Summarizing, a unimodular ribbon Hopf algebra yields an invariant $J_{H}$ of $2$-handlebodies through a rather complicated process. Fortunately, we do not have to go through the details since a direct algorithm computing the invariant is given in [BdR23, Section 8]. Let us sketch the algorithm, since it will have an importance in the proofs below.

### v2

@@ 1 Introduction
Unimodular, ribbon categories introduced by Turaev in [Tur94] constitute an important topic in quantum algebra, chiefly for their many applications in quantum topology. Since the 1990s, multiple constructions of 3-manifold and 4-manifold invariants have been constructed using these essential ingredients [Tur94], [KL01], [CY93]. Ribbon Hopf algebras are a natural source of such categories through their representation theory. A ribbon Hopf algebra $H$ is a quasitriangular Hopf algebra (i.e. it comes with the extra data of an invertible element $R$ of $H\otimes H$ called the R-matrix) together with a central element of $H$ called the ribbon element. Both $R$ and $v$ are required to satisfy certain equations, see Section 2.
In particular, one may define a $3$-manifold invariant from any so-called factorizable unimodular ribbon Hopf algebra. Factorizability is a non-degeneracy condition related to the monodromy matrix $M:=R_{21}R$ of the algebra, where $R_{21}$ is the R-matrix with permuted entries (see Proposition 2.10). Recall also that a Hopf algebra $H$ is unimodular if it admits an integral $\lambda\in H^{*}$ and a cointegral $\Lambda\in H$ satisfying certain properties, as detailed in Definition 2.5.
@@
At the topological level, non-factorizability is directly related to the ability of the invariants from [BdR23] to differentiate between $1$-handles and $2$-handles, which in turn is related to the ability of the invariant to see beyond the boundary of the $4$-dimensional $2$-handlebody. Non-(co)semisimplicity is related to the sensitivity of the invariant to a stabilization operation on handlebodies. This is detailed in Section 5.
@@ 1.1 Main results
Hence, in an attempt to avoid such “degeneration to the boundary”, in this paper, we introduce the stricter notion of strong non-factorizability. We demand that the result of applying the integral to either side of the monodromy matrix, with the antipode previously applied to its first tensor component, is zero, and hence cannot be related to the cointegral by any invertible element of the algebra:
$\big{(}(\lambda\circ S)\otimes\mathrm{Id}\big{)}M=\big{(}S\otimes\lambda\big{)}M=0.$
@@
First we discuss a family of Hopf algebras that was introduced in [BDG00] and analysed from the angle of quasitriangularity and ribbonness in [Nen04]. Thus, we will call these Hopf algebras Nenciu algebras. Let $\mathbf{m}\in\mathbb{Z}^{s}$ be a row tuple of integers of length $s\in\mathbb{Z}$, $t\in\mathbb{Z}$ and $\boldsymbol{\mathsf{d}},\boldsymbol{\mathsf{u}}$ be $t\times s$ matrices with integer entries. A Nenciu algebra is generated by grouplike generators $K_{a}$, $a=1,\dots,s$ such that $K^{m_{a}}_{a}=\mathbf{1}$, commuting with each other, as well as skew-primitive nilpotent generators $X_{k}$, $k=1,\dots,t$ such that $X^{2}_{k}=0$. The relations between the generators are prescribed by the matrices $\boldsymbol{\mathsf{d}},\boldsymbol{\mathsf{u}}$ and are diagonal, that is commuting any two elements results only in a constant in front of the swapped product. We denote the resulting algebra by $H(\mathbf{m},t,\boldsymbol{\mathsf{d}},\boldsymbol{\mathsf{u}})$, and a full definition is provided in Definition 3.1. In [Nen04], an array of previously existing quasitriangular Hopf algebra constructions, including the family $\operatorname{SF}_{2n}$, were unified, and sufficient conditions for such an algebra to be ribbon where determined, but no new examples were given. In particular, the questions of unimodularity and factorizability were not adressed.
In this paper, we exhibit new examples. Moreover we determine sufficient conditions for unimodularity, and for a number of instances we establish when non-factorizable and strongly non-factorizable ribbon structures are admitted, relying in particular on a theorem of Radford [Rad94], reproduced below as Theorem 2.9. We also study another property: the algebra is called anomaly-free if $\lambda(v)$ is invertible and anomalous otherwise (see Definition 2.20). Topologically, as explained in [BdR23], when the algebra is both factorizable and anomaly-free, the $4$-dimensional $2$-handlebody invariant associated to the algebra is actually an invariant of the boundary of the handlebody. For completeness, we check that anomaly-freeness is not satisifed in the examples where the corresponding quasitriangular structure is strongly non-factorizable. Combining the results of [Nen04] with our own, we have the following.
@@
However, we again observed that no true 4-dimensional information seems to be carried by the $4$-dimensional $2$-handlebody invariants given by this family of Hopf algebras. More precisely, computations for multiple examples suggest that the invariant might be given by the product of the 2-dimensional invariant of the spine of the handlebody obtained from $H(\mathbf{m},t,\boldsymbol{\mathsf{d}},\mathbf{u})$ and the refined 3-dimensional invariant of its boundary associated to $u_{q}\mathfrak{sl}_{2}$. Thus, we leave it as Conjecture 5.11.
@@ 1.2 Structure of the paper
The paper is organised as follows: in Section 2 we introduce all the relevant definitions pertaining to unimodular ribbon Hopf algebras and the notation used to describe the Nenciu algebras. In Section 3 the Nenciu algebras are introduced together with the conditions for a given one to admit unimodular, ribbon, and strongly non-factorizable structures. In Section 4, we introduce the second family of examples and once again give conditions for strong non-factorizability. In Section 5 we recall some topological background, including the Kerler-Lyubashenkno invariants of 4-dimensional 2-handlebodies, and discuss the efficacy of the invariants provided by our examples. Finally, in Appendix A we list all the examples of Hopf algebras found throughout the paper, and in Appendix B we collect some proofs that are too long to be presented in the main body.
Acknowledgements: The authors would like to thank Ivelina Bobtcheva, Marco de Renzi, Azat Gainutdinov, Cris Negron and Matthieu Faitg for early comments on the paper and fruitful discussions, as well as Riccardo Piergallini for making his Mathematica computer program available, and Anna Beliakova for her help and introducing them to the topic. Both authors were supported by Simons Collaboration on New Structures in Low-Dimensional Topology and Grant 200020_207374 of the Swiss National Science Foundation.
@@ 2.3 Quasitriangular and ribbon structures
A unimodular, quasitriangular Hopf algebra $(H,R)$ with an integral $\lambda\in H^{*}$ is called strongly non-factorizable if any of the two equivalent conditions is fulfilled
$\lambda(S(M^{\prime}))M^{\prime\prime}=0,$
@@
$S(M^{\prime})\lambda(M^{\prime\prime})=0.$
@@ 3 The Nenciu construction
The Nenciu construction of [Nen04] produces finite-dimensional ribbon Hopf algebras generated by grouplike generators collected in a tuple $\mathbf{K}=(K_{1},\dots,K_{s})$ and nilpotent generators collected in a tuple $\mathbf{X}=(X_{1},\dots,X_{t})$, such that the group subalgebra $\mathbb{C}[\mathbf{K}]$ is abelian, the nilpotent generators are skew-primitive with respect to $\mathbb{C}[\mathbf{K}]$ and all commutation relations between generators are diagonal. The choice of the generators and relations is strongly constrained, and sufficient conditions for the result to be a ribbon Hopf algebra were determined in [Nen04]. We will see that strongly non-factorizable examples of this construction exist.
@@ 3.1 Hopf algebra structure
$\boldsymbol{\xi}^{\boldsymbol{\mathsf{d}}_{k}\cdot\boldsymbol{\mathsf{u}}_{k}}=-1$
@@ 3.4 Quasitriangular structure
The following is a Corollary of Proposition 3.16.
@@
    since $\boldsymbol{\mathsf{d}}_{k}\boldsymbol{\mathsf{z}}=-\boldsymbol{\mathsf{u}}_{k}$. In the latter expression combining the two sums yields
@@
    Each piece of the rightmost sum corresponds in the former tensor factor to a monomial in $X_{k}$ generators, $\mathbf{X}^{\mathbf{r}}$, and in the latter tensor factor - its image under the composition of the duality (which, recall, is only a linear map), and the bialgebra map so the element $f((\mathbf{X}^{\mathbf{r}})^{*})\in H^{cop}$. The parameterization coefficients $\alpha_{kl}$ control whether a term appears in the sum, or equivalently in the image of $f$, or not, and with which coefficient.
@@ 3.6 Examples revisited
Here we stick to the notation for the nilpotent generators from the previous examples emphasizing the types. In particular, let $Z^{+}_{l}$ and $Z^{-}_{l}$ be nilpotent generators in $H(\mathbf{m},t,\boldsymbol{\mathsf{d}},\boldsymbol{\mathsf{u}})$ of opposite type, with neighbouring indices, that is $Z^{+}_{l}=X_{k}$ and $Z^{-}_{l}=X_{k+1}$ for some $k$. Let moreover $\{Z_{l}^{\pm},X_{k}\}=0$ for all $X_{k}$, including $Z_{l}^{\pm}$ themselves. We also made a change in the notation in the expression above and we denoted $\alpha_{k,k+1}=-\alpha_{k+1,k}:=\alpha_{l}$. We collect $\alpha_{l}$ in the tuple $\boldsymbol{\alpha}$.
@@
Finally, from the definitions of the integral and the R-matrix, we observe that all nilpotent generators except $Z^{\pm}_{l}$ (or all in Example 3.10) are absent from the respective R-matrices. There is no relation available that introduces them, as the commutations are necessarily diagonal, thus they are absent from the monodromy matrix as well. Moreover, the antipode does not change the content of nilpotent generators in any monomial. Since the integral is non-zero only on the linear combinations of monomials containing the top element $T=\mathbf{X}^{(1,\dots,1)}$, and since $T$ contains all the nilpotent generators, we retrieve the result. ∎
@@ 4.1 The small quantum group $u_{q}\mathfrak{sl}_{2}$
We use this section to set the conventions for the small quantum group. We follow mostly the conventions of [CP95, Section 9.1] and [EGNO16, Section 5.7], used also in [BdR23]. Let $r^{\prime}=r/2$ and $r^{\prime\prime}=r/4$ throughout.
@@ 4.2 The semi-direct biproduct structure
We emphasize that all tensor products between $U$ and $H$ in Definition 4.4 are taken in the category $\text{{Vect}}_{\mathbb{C}}$, which carries the trivial braiding. In particular, this is a different setting than the Radford biproduct which is taken in the braided category of Yetter-Drinfeld modules over the acting algebra, see [AS06].
@@ 4.3 Non-factorizable biproducts with $u_{q}\mathfrak{sl}_{2}$
This follows directly from the fact that both $U$ and $H$ have monomial bases, by the Poincaré-Birkhoff-Witt theorem (see for instance [Kas12]) and Proposition 3.6, respectively. Indeed, all newly introduced relations between the two bases are diagonal by Theorem 4.7. Hence, by the same argument as in the proof of Proposition 3.6, we can establish the existence of the monomial basis. ∎
@@
    and the corresponding pivotal element $g$ and the Drinfeld element
    $u_{H}:=g^{-1}\exp\left(-2\sum^{t_{2}}_{k=1}\alpha_{l}Z^{+}_{l}Z^{-}_{l}\right).$
@@
By the reasoning of Proposition 3.28, if $H=H(\mathbf{m},t,\boldsymbol{\mathsf{d}},\boldsymbol{\mathsf{u}})$ is strongly non-factorizable, the top element $\mathbf{X}^{(1,\dots,1)}$ is missing from the monodromy matrix $M_{H}\in H\otimes H$. Since all new relations are also diagonal, and as before the antipode does not change the content of nilpotent generators in any monomial, it will be missing from the new monodromy matrix $M\in(U\ltimes H)\otimes(U\ltimes H)$, so by the same argument the result follows. ∎
@@ 4.4 Examples
$\Delta(K_{1})=K_{1}\otimes K_{1},\;\;\Delta(K_{2})=K_{2}\otimes K_{2},\;\;\Delta(X^{\pm})=\mathbf{1}\otimes X^{\pm}+X^{\pm}\otimes L^{\pm 1}$
@@
$\Delta(K_{1})=K_{1}\otimes K_{1},\;\;\Delta(K_{2})=K_{2}\otimes K_{2},\;\;\Delta(X^{\pm})=\mathbf{1}\otimes X^{\pm}+X^{\pm}\otimes\bar{L}^{\pm 1}$
@@
    $\Delta(E)=\mathbf{1}\otimes E+E\otimes K$
@@
Thus, we see the top monomial picked up by the integral of $U\ltimes H$ is absent from the monodromy matrix, as in Proposition 3.28. We conclude the proof with the remark that we need not ask about the fate of generators $K_{a}$ under the map $f$, as this does not affect the strong non-factorizability condition in this case. ∎
@@ 5.2 Invariants of handlebodies up to $2$-equivalence
A very important feature of a unimodular ribbon Hopf algebra $H$ is that it can be “transmuted” into a $\operatorname{BP}$-Hopf algebra $\underline{H}$ in the category $H$-$\operatorname{mod}$ (as a special case of [BdR23, Prop. 7.3]), where $\underline{H}$ is $H$ as an $H$-module for the left adjoint action from Proposition 3.8. Summarizing, a unimodular ribbon Hopf algebra yields an invariant $J_{H}$ of $2$-handlebodies through a rather complicated process. Fortunately, we do not have to go through the details since a direct algorithm computing the invariant is given in [BdR23, Section 8]. Let us sketch the algorithm, since it will have an importance in the proofs below.